OpenStudy (anonymous):

x+4=3[x-2(x+1)] A -10 B -5/2 C 5/2 D 10 I keep getting 1 which aren't answer choices.

OpenStudy (anonymous):

show your work, please

OpenStudy (anonymous):

I simplified to -2x=-2

OpenStudy (anonymous):

x+4=3[x-2(x+1)] x+4=3[x-2x+2] x+4=3x-6x+6 x-2=3x-6x x-2=-3x -2=-2x

OpenStudy (anonymous):

@curiouscrab post the whole stuff, please.

OpenStudy (anonymous):

the second line: